# A Weight-aware-based Multi-source Unsupervised Domain Adaptation Method for Human Motion Intention Recognition

### v1

@@ V-B Feature Alignment and Classification Results
The Classification Results on ENABL3S and DSADS: To answer $\mathbf{Q_{1}}$, we compare different domain adaptation methods in ENABL3S and DSADS datasets, including DANN [20], MMD [21], MCD [22], DFA [55], GFA [16] and EDHKD [17]. The classification results on ENABL3S and DSADS of these methods come from article [17], where EDHKD is the state-of-the-art method for HMI using domain adaptation technology. Table II The detailed results for different target subjects on ENABL3S and DSADS datasets. Table III compares the mean accuracy of classification on the target subject for different domain adaptation methods. The test time of methods on this table is close to 1 ms, which guarantees the real-time of the test.
Table III shows that WMDD achieves the accuracy of $94.6\%\pm 2.0\%$ and $99.0\%\pm 0.7\%$ for target subject in ENABL3S and DSADS datasets respectively, which are $0.2\%$ and $1.6\%$ higher than the state-of-the-art HMI result using domain adaptation methods. This verifies that WMDD can achieve a better performance than previous methods in both ENABL3S and DSADS datasets. However, the improvement in ENABL3S is not significant despite that the samples of ENABL3S are larger than DSADS. The reason lies in the effects of transition activities. This results of classification are consistent with that of feature alignment.
V-C Ablation Experiments
Effects of Weight $\bm{\alpha}$ and Adversarial Learning: To answer $\mathbf{Q_{2}}$, the ablation experiments for two parts are designed (see Table IV). The weight $\bm{\alpha}$ reflects the discrepancy between different source domains and target domain. The adversarial learning is used to improve the classifier’s performance. From this table, the performance in ENABL3S and DSADS is improved by $1.07\%$ and $1.02\%$ respectively when considering the difference between source domains. The performance is improved by $1.72\%$ and $7.25\%$ respectively when introducing the adversarial learning between the feature generator and classifier. Therefore, the usage of domain weight and adversarial learning can both improve the performance of the classifier, and the effect of adversarial learning is more significant.
Figure 5: (a) The relationship between classification accuracy and the number of classifiers $K$ on ENABL3S and DSADS datasets. The red line and blue line represent the classification accuracy of ENABL3S and DSADS datasets respectively. (b) The relationship between classification accuracy and trade-off coefficient $\eta$ on ENABL3S and DSADS datasets. The red dotted line represents the relatively optimal parameters on two datasets.
@@ V-C Ablation Experiments
To answer $\mathbf{Q_{3}}$, we give the training curve and MDD under margin coefficient $\gamma=0.01,0.1,1$ on ENABL3S and DSADS datasets. Fig. 4 shows that the method WMDD achieves better classification accuracy under margin coefficient $\gamma=0.1$ for ENABL3S and DSADS datasets. The recognition accuracy are $91.6\%$, $94.6\%$, $92.0\%$ under $\gamma=0.01$, $0.1$, $1$ respectively for ENABL3S dataset, and are $97.9\%$, $99.0\%$, $97.5\%$ for DSADS dataset. The final average MDD values are $0.06$, $0.34$, $2.20$ under $\gamma=0.01$, $0.1$, $1$ respectively for ENABL3S dataset, and are $0.06$, $0.35$, $3.84$ for DSADS dataset. The results indicate that the average MDD between each source subject and target subject increases as margin coefficient $\gamma$ increases. However, for too large $\gamma$, the recognition accuracy cannot be improved. This shows that the margin $\mu$ in Theorem 3 cannot be set too large.
Effects of the Number of Classifiers $K$ and Trade-off Coefficient $\eta$: To answer $\mathbf{Q_{3}}$, the more experiments for the number of classifiers $K$ and trade-off coefficient $\eta$ are conducted below. The more classifiers often consume more computation cost. It is a necessity to choose an appropriate number of classifiers to achieve a balance between the classification performance and computation cost. The coefficient $\eta$ is used to trade off the fitting ability and generalization ability of the network.
Fig. 5 (a) gives the relationship between classification accuracy and the number of classifiers $K$ on ENABL3S and DSADS datasets. This shows that it’s not that the more classifiers there are, the better the classification performance. Compared with the accuracy of $5$ classifiers, the accuracy of $25$ classifiers decreases by $1.56\%$ and $5.61\%$ for ENABL3S and DSADS datasets respectively. For lightweight networks, only by selecting an appropriate number of classifiers can classification performance be effectively improved. Fig. 5 (b) shows the relationship between classification accuracy and trade-off coefficient $\eta$ on ENABL3S and DSADS datasets. The results indicate that too large or too small trade-off coefficient $\eta$ will decrease the performance of the algorithm. The relatively optimal parameters for $K$ and $\eta$ are both $5$ on ENABL3S and DSADS datasets.
@@ VI Discussion
This paper aims to extend MDD theory to multi-source UDA, and propose a novel multi-source UDA algorithm for HMI recognition based on the derived theory. In the below part, we give further discussions for multi-source UDA theory and algorithm.
@@ VI-A The Discussion for Theory
Challenge 2: The dataset is collected from multiple source domains that might be different not only from the target domain but also from each other. The source domain weight determined by MDD is incorporated into Theorem 1 to measure the difference between source domains, where the source domain with the small gap to the target domain should have a large weight. The weight is adjusted adaptively by MDD according to Eq. (14). Since the training data is randomly selected, the drawn data hardly fully describe the entire distribution characteristic. Adjusting weight dynamically according to the drawn data is beneficial to improving performance on the target domain. Table IV also verifies this conclusion.
@@ VI-B The Discussion for Algorithm
Since the margin loss easily leads to gradient vanishing, the cross-entropy loss is employed to replace the margin loss in this paper. The $\gamma$ in Eq. (18) reflects the margin $\mu$ in MDD. Theorem 4 shows that the result of using a surrogate function can better measure the discrepancy between different distributions and maintain the main characteristics of MDD through $\gamma$. The algorithm is essentially a two-stage game problem. The first stage is stackelberg game, and the second stage is zero-sum game. The details are given below.
1) Stackelberg game: Eq. (12) can be regarded as stackelberg game model, where auxiliary classifiers $\bm{f^{\prime}}_{j}$ is set as leader and classifiers $\bm{f}$ is follower. In this stage, the auxiliary classifiers are first updated and the MDD between source and target domain is estimated. Then classifiers and feature generator are updated together through minimizing the generalization error. 2) Zero-sum game: Eq. (22) can be regarded as two-player zero-sum game model, where feature generator $\phi$ and classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$ are the adversarial parties. The generalization ability (performance on target domain) is improved through constant adversarial learning between feature generator and classifiers.
@@ VII Conclusion
This paper developed a new theory for multi-source UDA based on margin disparity discrepancy and derived a novel generalization bound for multi-source UDA. Motivated by generalization bound, a novel weight-aware-based multi-source UDA algorithm (WMDD) was proposed for HMI recognition. The proposed method WMDD can improve classification accuracy of HMI recognition tasks through considering the difference between each source subject, where the source domain weight can be adjusted adaptively by the estimated discrepancy. WMDD can also guarantee the real-time of HMI recognition by utilizing a lightweight network. Extensive experiments confirmed that WMDD can achieve state-of-the-art accuracy on HMI recognition tasks. We expect the proposed theory and algorithm can provide reference and inspiration for other multi-source UDA theories and cross-subject application fields.
-A Proof of Theorem 1
@@
where the second inequality is the important property of margin loss for any $\mu$ and $\bm{f}$, and $\lambda=\mathcal{L}_{Q}^{\mu}\left(\bm{f^{*}}\right)+\mathcal{L}_{P}^{\mu}\left(\bm{f^{*}}\right)=\min_{\bm{f}\in\mathcal{F}}\{\mathcal{L}_{Q}^{\mu}\left(\bm{f}\right)+\mathcal{L}_{P}^{\mu}\left(\bm{f}\right)\}$. Therefore, the proof of Theorem 1 is completed.
-B Proof of Theorem 2
@@
Let $\mathcal{G}$ be a family of margin loss functions defined as $\mathbb{E}_{\bm{x}\sim\mathcal{D}_{x}}\big{\{}\mathbb{W}_{\mu}\big{[}\omega_{\bm{f^{\prime}}}\big{(}x,h_{\bm{f}}(x)\big{)}\big{]}\big{\}}$ and $\Omega_{2}(\mathcal{F})$ be a family of functions defined in Definition 2. Then, for any sample $\widehat{\mathcal{D}}_{x}=\{(\bm{x}_{i}\}_{i=1}^{m}$, the following relation holds between the empirical Rademacher complexities of $\mathcal{G}$ and $\Omega_{2}(\mathcal{F})$:
@@
where $\widehat{\mathcal{D}}_{s}$ and $\widehat{\mathcal{D}}$ are datasets sampled from the mixture distribution $\sum_{j=1}^{N}\alpha_{j}P_{j}$ and $Q+\sum_{j=1}^{N}\alpha_{j}P_{j}$, respectively. Therefore, the proof of Theorem 3 is completed.
-D Proof of Theorem 4
@@
$=$ $\int_{\bm{x}\in\mathcal{X}}\gamma P(\bm{x})\log\Big{[}\frac{\gamma P(\bm{x})}{\frac{\gamma P(\bm{x})+Q(\bm{x})}{\gamma+1}}\Big{]}\mathrm{d}x$

### v2

@@ V-C Ablation Experiments
To answer $\mathbf{Q_{3}}$, we give the training curve and MDD under margin coefficient $\gamma=0.01,0.1,1$ on ENABL3S and DSADS datasets. Fig. 5 shows that the method WMDD achieves better classification accuracy under margin coefficient $\gamma=0.1$ for ENABL3S and DSADS datasets. The recognition accuracy are $91.6\%$, $94.6\%$, $92.0\%$ under $\gamma=0.01$, $0.1$, $1$ respectively for ENABL3S dataset, and are $97.9\%$, $99.0\%$, $97.5\%$ for DSADS dataset. The average values of the estimated MDD are $0.06$, $0.34$, $2.20$ under $\gamma=0.01$, $0.1$, $1$ respectively for ENABL3S dataset, and are $0.06$, $0.35$, $3.84$ for DSADS dataset. This result indicates that the average MDD between each source subject and target subject increases as margin coefficient $\gamma$ increases. However, for too large $\gamma$, the recognition accuracy cannot be improved. This shows that the margin $\mu$ in Theorem 3 cannot be set too large.
Effects of the Number of Classifiers $K$ and Trade-off Coefficient $\eta$: To answer $\mathbf{Q_{3}}$, the more experiments for the number of classifiers $K$ and trade-off coefficient $\eta$ are conducted below. The more number of classifiers $K$ often consume more computation cost. It is a necessity to choose an appropriate number of classifiers $K$ to achieve a balance between the classification performance and computation cost. The coefficient $\eta$ is used to trade off the fitting ability and generalization ability of the network. The performance of WMDD is determined by this hyperparameter.
Table IV shows the relationship between the classification accuracy and the number of classifiers $K$, and the relationship between classification accuracy and trade-off coefficient $\eta$ on ENABL3S and DSADS datasets. This suggests that the increase in the number of classifiers $K$ does not necessarily lead to improved classification performance. Compared with the accuracy of $5$ classifiers, the accuracy of $25$ classifiers decreases by $1.56\%$ and $5.61\%$ for ENABL3S and DSADS datasets respectively. For lightweight networks, classification performance can be effectively enhanced by selecting an optimal number of classifiers. Moreover, the results indicate that the relatively optimal parameters for $K$ and $\eta$ are both $5$ on ENABL3S and DSADS datasets. If coefficient $\eta$ is too small, the network’s fitting ability will improve, but its generalization capability will decline. Conversely, if coefficient is too large, the generalization ability will increase, while its fitting ability will diminish. Therefore, too large or too small coefficient $\eta$ will decrease the performance of the algorithm.
@@ VI Discussion
This paper aims to extend MDD theory to multi-source UDA, and propose a novel multi-source UDA algorithm for HMI recognition based on the derived theory. In the below part, we give further discussions for multi-source UDA theory and algorithm, and discuss more related works about unsupervised domain adaption.
@@ VI-A The Discussion for Theory
Challenge 2: The dataset is collected from multiple source domains that might be different not only from the target domain but also from each other. The source domain weight determined by MDD is incorporated into Theorem 1 to measure the difference between source domains, where the source domain with the small gap to the target domain should have a large weight. The weight is adjusted adaptively by MDD according to Eq. (13). Since the training data is randomly selected, the drawn data hardly fully describe the entire distribution characteristic. Adjusting weight dynamically according to the drawn data is beneficial to improving performance on the target domain. Table V also verifies this conclusion.
@@ VI-B The Discussion for Algorithm
Since the margin loss easily leads to gradient vanishing, the cross-entropy loss is employed to replace the margin loss in this paper. The $\gamma$ in Eq. (17) reflects the margin $\mu$ in MDD. Theorem 4 shows that the result of using a surrogate function can better measure the discrepancy between different distributions and maintain the main characteristics of MDD through $\gamma$. The algorithm is essentially a two-stage game problem. The first stage is the game problem between the auxiliary classifiers and classifiers, and the second stage is the game problem between the classifiers and feature generator. In the first stage, the auxiliary classifiers are first updated and the MDD between source and target domain is estimated. Then classifiers and feature generator are updated together through minimizing the generalization error. In the second stage, the feature generator and classifiers are the adversarial parties. The generalization ability (performance on target domain) is improved through constant adversarial learning between feature generator and classifiers.
VI-C The Discussion for Unsupervised Domain Adaption
Domain shift is a fundamental challenge for recognition and prediction, where there is a gap between the training (source) domain and test (target) domain. Domain adaption method and meta learning technique have been utilized to deal with the above problem. Meta learning, such as MAML [50], acquires knowledge from multiple tasks, enabling faster adaptation and generalization to new tasks. Domain adaption can be considered as part of meta learning when treating domain as a particular task. The most unsupervised domain adaption methods [12, 11], including our method, aim to achieve feature alignment between source and target domain. However, the performance of these methods would be limited if data in source domain is insufficient [51]. Some meta learning techniques [52] can be combined with domain adaptation to solve the above problem. The goal of domain adaption and meta learning is same, that is pursuing the generality and adaptability in artificial intelligence.
Moreover, unsupervised domain adaption technique has been widely applied in the cross-subject tasks. Some methods make improvements to the perception network framework itself, such as using multi-encoder autoencoder [65], to improve performance across subjects. Other methods improve the algorithm itself, such as modifying the loss function [32, 58], to minimize the error in the target subject and thus improve the recognition accuracy. Compared with the above cross-subject methods, the proposed method of this paper can fully integrate the differences between multiple subject data through adaptively adjusting the weight of different source subjects, and then the more effective information can be obtained for classification. Therefore, under the same amount of data, the recognition accuracy of the proposed method can be improved on the cross-subject tasks.
@@ VII Conclusion
This paper developed a new theory for multi-source UDA based on margin disparity discrepancy and derived a novel generalization bound for multi-source UDA. Motivated by generalization bound, a novel weight-aware-based multi-source UDA algorithm (WMDD) was proposed for HMI recognition. The proposed method WMDD can improve classification accuracy of HMI recognition tasks through considering the difference between each source subject, where the source domain weight can be adjusted adaptively by the estimated discrepancy. WMDD can also guarantee the real-time of HMI recognition by utilizing a lightweight network. Extensive experiments confirmed that WMDD can achieve state-of-the-art accuracy on HMI recognition tasks.
We expect the proposed theory and algorithm can provide reference and inspiration for other multi-source UDA theories and cross-subject application fields. However, the current domain adaptation methods require retraining the network for each new target subject to achieve better performance, which is computationally expensive. Future works can focus on exploring a domain-adaptive fine-tuning method that leverages unlabeled data from the target subject to fine-tune the base model, thereby enhancing performance.
Supplementary Materials (Appendix)
VII-A Proof of Theorem 1
@@
where the second inequality is the important property of margin loss for any $\mu$ and $\bm{f}$, and $\lambda=\mathcal{L}_{Q}^{\mu}\left(\bm{f^{*}}\right)+\mathcal{L}_{P}^{\mu}\left(\bm{f^{*}}\right)=\min_{\bm{f}\in\mathcal{F}}\{\mathcal{L}_{Q}^{\mu}\left(\bm{f}\right)+\mathcal{L}_{P}^{\mu}\left(\bm{f}\right)\}$. This completes the proof of Theorem 1.
VII-B Proof of Theorem 2
@@ VII-C Proof of Theorem 3
Let $\mathcal{G}$ be a family of margin loss functions defined as $\mathbb{E}_{\bm{x}\sim\mathcal{D}_{x}}\big{\{}\mathbb{W}_{\mu}\big{[}\omega_{\bm{f^{\prime}}}\big{(}x,h_{\bm{f}}(x)\big{)}\big{]}\big{\}}$ and $\Omega_{2}(\mathcal{F})$ be a family of functions defined in Definition 2. Then, for any sample $\widehat{\mathcal{D}}_{x}=\{\bm{x}_{i}\}_{i=1}^{m}$, the following relation holds between the empirical Rademacher complexities of $\mathcal{G}$ and $\Omega_{2}(\mathcal{F})$:
@@
where $\widehat{\mathcal{D}}_{s}$ and $\widehat{\mathcal{D}}$ are datasets sampled from the mixture distribution $\sum_{j=1}^{N}\alpha_{j}P_{j}$ and $Q+\sum_{j=1}^{N}\alpha_{j}P_{j}$, respectively. This completes the proof of Theorem 3.
VII-D Proof of Theorem 4
@@
$=$ $\int_{\bm{x}\in\mathcal{X}}\gamma P(\bm{x})\left\{\log\Big{[}\frac{\gamma P(\bm{x})}{\frac{\gamma P(\bm{x})+Q(\bm{x})}{\gamma+1}}\Big{]}-\log\left[\frac{\gamma+1}{\gamma}\right]\right\}\mathrm{d}x$
$+$ $\int_{\bm{x}\in\mathcal{X}}Q(\bm{x})\left\{\log\Big{[}\frac{Q(\bm{x})}{\frac{P(\bm{x})+Q(\bm{x})}{\gamma+1}}\Big{]}-\log\left[\gamma+1\right]\right\}\mathrm{d}x$
$=$ $\int_{\bm{x}\in\mathcal{X}}\gamma P(\bm{x})\log\Big{[}\frac{P(\bm{x})}{\frac{\gamma P(\bm{x})+Q(\bm{x})}{\gamma+1}}\Big{]}\mathrm{d}x$